I am not sure if this is the correct place for this and I do not understand it at all. Can you help please. Thanks so much for your time.
Evaluate 66 ft/s times 1 mi/5,280 ft times 3,600 s/h.

Answer by edjones(7311)

66ft=66/5280=1/80 mi
(1/80)/(1/3600)
1/80*3600=45 mph
Ed